Reports of 33080
PEM (Center)
takaaki.yokozawa - 8:59 Thursday 27 November 2025 (35685) Print this report
PEM injection 251127
I performed the PEM injection tests

Silent run - 5:12

1. Shake injection to TMSX table EXC2
11/27 05:42:00 - 06:17:00 (JST)
EXC : K1:PEM-EXCITATION_EX0_RACK_2_EXC
REF : K1:PEM-ACC_TMSX_TABLE_TMS_Z_OUT_DQ
info : TMSX EXC2 50 - 900 Hz 171 point 5 Hz resolution 50cnt excitation

2. Shake injection to TMSX table EXC3
11/27 06:17:00 - 06:47:00 (JST)
EXC : K1:PEM-EXCITATION_EX0_RACK_3_EXC
REF : K1:PEM-PORTABLE_EXC_RACK_EX0_ADC0_DSUB26_OUT_DQ
info : TMSX EXC3 50 - 900 Hz 171 point 5 Hz resolution 50cnt excitation

3. Shake injection to TMSX table EXC4
11/27 06:47:00 - 07:27:00 (JST)
EXC : K1:PEM-EXCITATION_EX0_RACK_4_EXC
REF : K1:PEM-PORTABLE_EXC_RACK_EX0_ADC0_DSUB27_OUT_DQ
info : TMSX EXC4 50 - 900 Hz 171 point 5 Hz resolution 100cnt excitation

Silent run 7:27 - 7:47

4. Acoustic injection to REFL
11/27 07:47:00 - 08:22:00 (JST)
EXC : K1:PEM-EXCITATION_MCF0_RACK_14_EXC
REF : K1:PEM-MIC_PR_BOOTH_REFL_Z_OUT_DQ
Info : REFL 50 - 900 Hz 171 point 5 Hz resolution 100 cnt excitation

5. Acoustic injection to OMC
11/27 08:22:00 - 08:57:00 (JST)
EXC : K1:PEM-EXCITATION_SR3_RACK_6_EXC
REF : K1:PEM-MIC_OMC_BOOTH_OMC_Z_OUT_DQ
Info : OMC 50 - 900 Hz 171 point 5 Hz resolution 100 cnt excitation
During the measurement, SRM was tripped
MIF (ITF Control)
takafumi.ushiba - 8:16 Thursday 27 November 2025 (35684) Print this report
Comment to Recovering the 'INCREASING LAS POWER' state by reducing the IMC loop gain (35679)

>To address this, I reduced the IMC CMS in1gain by 3 dB

Did you reduce IN2 gain as well?
To keep the relative gain between IMC loop and CARM loop, we also need to change IN2 gain if IN1 gain is changed.

MIF (Noise Budget)
Hiroki Fujimoto - 4:55 Thursday 27 November 2025 (35683) Print this report
Noise measurement of CARM RFPD

Abstract

We measured the sensing noise (dark noise and shot noise) of the CARM RFPD at REFL as part of the CARM noise budget.
By analyzing the power dependence of the demodulated noise spectra, we extracted the dark noise level and the shot noise coefficient, as well as the effective transimpedance of the RFPD.

Details

As part of the CARM noise budget, we measured the sensing noise (dark noise + shot noise) of the CARM RFPD at REFL.
The procedure was as follows: the reflected beam from the PRM was injected into the CARM RFPD, and the demodulated noise spectrum was measured for several PD input powers.
By fitting the power dependence of these spectra, we estimated the dark noise and the coefficients associated with shot noise.
The demodulated spectra were first amplified with a SR560 (high-pass @ 3 Hz, gain = 100) and then measured with a Moku:Lab.

Figure 1 shows the demodulated noise spectra taken at different PD input powers (the amplification by the SR560 has been corrected).
To extract the noise-floor values for each input power, we computed the mean and standard error over the frequency band from 50 kHz to 100 kHz for each spectrum.
This frequency region was selected to avoid the peak near 40 kHz.

Figure 2 shows the power dependence of the obtained noise-floor values.
The data were fitted using the following model to extract the power dependence of the PD dark noise and shot noise:

Fitting model:
δVRFPDnoise=δVdark2+δVshot2\delta V_\mathrm{RFPDnoise} = \sqrt{\delta V_\mathrm{dark}^2+\delta V_\mathrm{shot}^2}
Definitions:
δVdark\delta V_\mathrm{dark} : Dark noise of the RFPD
δVshotR2eα×Pin\delta V_\mathrm{shot} \equiv R\sqrt{2e\alpha}\times\sqrt{P_\mathrm{in}}: Shot noise after RFPD demodulation
RR: effective transimpedance including demodulation
ee: elementary charge
α\alpha : responsivity [A/W]
PinP_\mathrm{in}: input power to the RFPD

Result of fitting:
Vdark=3.96(1)×10-8V/HzV_\mathrm{dark}=3.96(1)\times10^{-8}\,\mathrm{V/\sqrt{Hz}}
R2eα=2.271(8)×10-8V/HzmW=7.18(2)×10-7V/HzWR\sqrt{2e\alpha}=2.271(8)\times10^{-8}\,\mathrm{V/\sqrt{Hz\,mW}} =7.18(2)\times10^{-7}\,\mathrm{V/\sqrt{Hz\,W}}

Assuming α=0.77A/W\alpha=0.77\,\mathrm{A/W} (QE = 0.9), the effective transimpedance RR and the power-to-voltage conversion efficiency including demodulation RαR\alpha are
R=1.447(5)×103ΩR=1.447(5)\times10^3\,\mathrm{\Omega}
Rα=1.114(4)×103V/WR\alpha=1.114(4)\times10^3\,\mathrm{V/W}

These results will be used for the CARM noise budget soon.

Images attached to this report
MIF (Noise Budget)
kentaro.komori - 3:36 Thursday 27 November 2025 (35682) Print this report
Measurement of transfer function and dark noise of CARM CMS

[Fujimoto, Tanaka, Komori]

Abstract:

We measured the transfer function and dark noise of the CARM CMS.
The results appear reasonable, and we have now completed the measurements necessary to perform the CARM noise budget.

Details:

We carried out the same set of measurements as those performed for the IMC CMS, as reported in klog:35669.
For the CARM CMS, we focus only on the fast path and ignore the slow path for now, because the slow path is used for MCE feedback and is negligible for the CARM noise budget.

The dark noise of the CARM CMS fast output was measured with a ×100 amplification using the SR560, and the result is shown by the blue line in Fig. 1.
We measured the spectrum in two frequency ranges, similar to yesterday’s IMC CMS measurements, and this time there was no discrepancy around the 1 kHz boundary.

The measured transfer function (blue dots) is compared with the modeled one (cyan line) in Fig. 2, and they agree well.
Using the modeled transfer function, I calculated the input-equivalent noise of the CARM CMS fast path (Fig. 3).

We also investigated the mismatch between spectra measured in two frequency bands reported in klog:35669.
Today, we found that one of the terminal resistors had a poor connection and was introducing excess noise. After replacing it, we successfully measured the IMC CMS slow-path dark noise with good agreement between the two spectral measurements (red in Fig. 1).

On the other hand, we observed a mismatch in the dark-noise measurement of the IMC CMS fast path (magenta in Fig. 1), even though it matched well in previous measurements.
We found that the dark noise is sensitive to mechanical stress applied to the CMS board connector.
Since the high-frequency measurement is reproducible and the IMC CMS fast-path dark noise is negligible at low frequencies, we will use the previous result.

The mismatch in the Mokulab dark noise is likely due to a degraded noise floor in the low-frequency band caused by additional downsampling.
Furthermore, we found that the overall gain discrepancy between measured and modeled transfer functions originated from the impedance setting of Mokulab: we should have divided the output by two. With this correction, the measured and modeled gains now match well.

We have completed the necessary measurements for the CARM noise budget and will proceed to the calculation phase.

Images attached to this report
MIF (ITF Control)
takafumi.ushiba - 0:24 Thursday 27 November 2025 (35681) Print this report
Comment to Recovering the 'INCREASING LAS POWER' state by reducing the IMC loop gain (35679)

As discussed in klog35456 and klog35472, high CARM gain might introduce the instability during increasing laser power.
So, 40kHz UGF would be good at least for achieving the stable switching of CARM input gains.

During the work reported in klog35472, I set CARM UGF as 40kHz, so the initially measured UGF seems unexpectedly too high.
I'm not so sure why the CARM UGF (and also IMC UGF) became so high, so it would be nice to investigate the reason.

MIF (ITF Control)
kenta.tanaka - 22:04 Wednesday 26 November 2025 (35680) Print this report
Comment to Recovering the 'INCREASING LAS POWER' state by reducing the IMC loop gain (35679)

This afternoon, we found that the lock loss was occured whenever the IN1GAIN value was changed from -23 dB to -22 dB after increasing the laser power itself. This process is performed after increasing the IMC output power. 

At first, we measured the CARM OLTF just before changing the gain values related the CARM loop (K1:IMC-SERVO_{IN1,FAST}GAIN, K1:LSC-REFL_SERVO_SLOW_GAIN). In this state, the OLTF itself should be the same as the state after passing through the INCRASING_LAS_POWER state. Fig. 1 shows the CARM OLTF, dark curve is the one before the adjustment. The UGF at that time is ~60 kHz, and the phase margin seems to be roughly 30 degrees. The gain changing process is performed to keep this OLTF so we considered that the phase margin is not enough to perform the changing process. Therefore, we attempted the process in the state which the overall gain reduced -3 dB by decreasing K1:IMC-SERVO_FASTGAIN, K1:LSC-REFL_SERVO_SLOW_GAIN to -3dB in advance. That is, the process started when IN1GAIN was -24dB, FASTGAIN was 1 dB, and SLOW_GAIN was 0.32

After changing the process, we succeeded in transiting the INCREASING_LAS_POWER state manually. The bright curve is the one after the adjustment. Current CARM UGF is ~40 kHz, phase margin is 43 deg. I wonder that the gain above ~ 150 kHz seems to be not changed even though we adjusted the overall gain. We need more investigation. 

Then, we modified the script in the INCREASING_LAS_POWER of the LSC_LOCK guardian. We tested the process by the guardian at twice, the first trial is failed on the way which FAST_GAIN decreased from 1 dB to  -2 dB. The second trial was succeeded. So we would like to know the succeess rate of this process. We left the IFO with this state.

  

Images attached to this comment
MIF (ITF Control)
kentaro.komori - 21:28 Wednesday 26 November 2025 (35679) Print this report
Recovering the 'INCREASING LAS POWER' state by reducing the IMC loop gain

[Fujimoto, Yokozawa, Ushiba (remote), Tanaka, Komori]

Abstract:

We resolved the lock-loss issue that occurred when increasing the laser power by tuning the IMC loop gain.

Details:

As reported in klog:35668, lock loss consistently occurred when the input power was increased.
Yokozawa-san found that the CARM open-loop gain exhibited an unexpected peak around 200 kHz (Fig. 1), which was a strong candidate for the cause of the lock loss.
Ushiba-san suggested that this could be due to an excessively high IMC open-loop gain.
I measured the IMC loop and confirmed that this was indeed the case: the UGF was around 200 kHz, and the phase margin was small—approximately 20 degrees or less (Fig. 2).

To address this, I reduced the IMC CMS in1gain by 3 dB, lowering the UGF to around 100 kHz and improving the phase margin to ~30 degrees (Fig. 3).
After this adjustment, we were able to increase the laser power without losing lock.

Tanaka-san updated the guardian setting for in1gain in the LOCK PREP state of the IMC guardian.
With this change, the system can now transition through the INCREASING LAS POWER state without lock loss.

Images attached to this report
Comments to this report:
kenta.tanaka - 22:04 Wednesday 26 November 2025 (35680) Print this report

This afternoon, we found that the lock loss was occured whenever the IN1GAIN value was changed from -23 dB to -22 dB after increasing the laser power itself. This process is performed after increasing the IMC output power. 

At first, we measured the CARM OLTF just before changing the gain values related the CARM loop (K1:IMC-SERVO_{IN1,FAST}GAIN, K1:LSC-REFL_SERVO_SLOW_GAIN). In this state, the OLTF itself should be the same as the state after passing through the INCRASING_LAS_POWER state. Fig. 1 shows the CARM OLTF, dark curve is the one before the adjustment. The UGF at that time is ~60 kHz, and the phase margin seems to be roughly 30 degrees. The gain changing process is performed to keep this OLTF so we considered that the phase margin is not enough to perform the changing process. Therefore, we attempted the process in the state which the overall gain reduced -3 dB by decreasing K1:IMC-SERVO_FASTGAIN, K1:LSC-REFL_SERVO_SLOW_GAIN to -3dB in advance. That is, the process started when IN1GAIN was -24dB, FASTGAIN was 1 dB, and SLOW_GAIN was 0.32

After changing the process, we succeeded in transiting the INCREASING_LAS_POWER state manually. The bright curve is the one after the adjustment. Current CARM UGF is ~40 kHz, phase margin is 43 deg. I wonder that the gain above ~ 150 kHz seems to be not changed even though we adjusted the overall gain. We need more investigation. 

Then, we modified the script in the INCREASING_LAS_POWER of the LSC_LOCK guardian. We tested the process by the guardian at twice, the first trial is failed on the way which FAST_GAIN decreased from 1 dB to  -2 dB. The second trial was succeeded. So we would like to know the succeess rate of this process. We left the IFO with this state.

  

Images attached to this comment
takafumi.ushiba - 0:24 Thursday 27 November 2025 (35681) Print this report

As discussed in klog35456 and klog35472, high CARM gain might introduce the instability during increasing laser power.
So, 40kHz UGF would be good at least for achieving the stable switching of CARM input gains.

During the work reported in klog35472, I set CARM UGF as 40kHz, so the initially measured UGF seems unexpectedly too high.
I'm not so sure why the CARM UGF (and also IMC UGF) became so high, so it would be nice to investigate the reason.

takafumi.ushiba - 8:16 Thursday 27 November 2025 (35684) Print this report

>To address this, I reduced the IMC CMS in1gain by 3 dB

Did you reduce IN2 gain as well?
To keep the relative gain between IMC loop and CARM loop, we also need to change IN2 gain if IN1 gain is changed.

PEM (Tool)
tatsuki.washimi - 16:16 Wednesday 26 November 2025 (35678) Print this report
Comment to Rain collector data at Atotsu looks strange (35667)

I checked the rain collector of the Atotsu weather station. It was fullfield by today's rain.

I and Sawada-san cleand it (removing algae).

Images attached to this comment
OBS (SDF)
dan.chen - 13:42 Wednesday 26 November 2025 (35677) Print this report
Comment to Changes of observation.snap during post O4c commissioning (35605)

We accepted the following SDF change caused by klog35676.

 

Images attached to this comment
CAL (XPcal)
dan.chen - 13:38 Wednesday 26 November 2025 (35676) Print this report
XPcal calibration

KAGRA Pcal-X updates (2025/11/26)

Workers: Dan Chen, Shingo Hido

We performed monthly Pcal-X calibration on 2025/11/26.

After the calibration, we updated EPICS parameters related to the Pcal-X system. No issues were found.

EPICS Key Before After Δ (After − Before)
K1:CAL-PCAL_EX_1_OE_R_SET 0.98527 0.98507 -0.00020
K1:CAL-PCAL_EX_1_OE_T_SET 0.98527 0.98507 -0.00020
K1:CAL-PCAL_EX_1_PD_BG_RX_V_SET -0.00383 -0.00381 0.00002
K1:CAL-PCAL_EX_1_PD_BG_TX_V_SET 0.00584 0.00572 -0.00012
K1:CAL-PCAL_EX_1_RX_V_R_SET 0.50205 0.50181 -0.00024
K1:CAL-PCAL_EX_2_INJ_V_GAIN 0.95254 0.95258 0.00004
K1:CAL-PCAL_EX_2_OE_R_SET 0.97694 0.97607 -0.00086
K1:CAL-PCAL_EX_2_OE_T_SET 0.97694 0.97607 -0.00086
K1:CAL-PCAL_EX_2_PD_BG_TX_V_SET 0.00462 0.00470 0.00009
K1:CAL-PCAL_EX_2_RX_V_R_SET 0.49795 0.49819 0.00024
K1:CAL-PCAL_EX_WSK_PER_RX_SET 1.48876 1.48709 -0.00167
K1:CAL-PCAL_EX_WSK_PER_TX1_SET 0.52745 0.52744 -0.00001
K1:CAL-PCAL_EX_WSK_PER_TX2_SET 0.38759 0.38765 0.00006

 

Images attached to this report
PEM (EX)
tatsuki.washimi - 12:57 Wednesday 26 November 2025 (35675) Print this report
Comment to PEM injection 251125 (35662)

I analyzed the PEM injection data:

2. Shake injection to TMSX table EXC1 (Fig.3.)
11/25 13:07:00 - 13:41:00 (JST)
EXC : K1:PEM-EXCITATION_EX0_RACK_1_EXC
REF : K1:PEM-POTABLE_EXC_RACK_EX0_ADC0_DSUB25_OUT_DQ
info : TMSX EXC1 50 - 900 Jz 171 point 5 Hz resolution 100cnt excitation

Note that the REF channel name was wrong.

Since a significant excess was not found for almost all frequency injection, the upper limit based on the Coupling Function model (w/o frequency-conversion) is also plotted. 

Images attached to this comment
AOS (General)
takahiro.yamamoto - 12:51 Wednesday 26 November 2025 (35674) Print this report
Comment to Strange vacuum pressure data; investigating the behavior of TMS-VIS during pumping down (35607)
I dumped the raw data of VAC-***_TMSX on 6/11, 2024 from the frame files
and confirmed that the data was surely filled as 0 before GPS=1402203200 (2024-06-12 13:53:02 JST).

Other channels at that time shows reasonable values, so it didn't seem to be a DAQ issue.
According to klog#29829, pumping down might be started before recovering CC-10.
PEM (EX)
tatsuki.washimi - 12:17 Wednesday 26 November 2025 (35673) Print this report
Comment to PEM injection 251125 (35662)

The channel name in the klog35662 was wrong

wrong K1:PEM-POTABLE_EXC_RACK_EX0_ADC0_DSUB25_OUT_DQ
correct K1:PEM-PORTABLE_EXC_RACK_EX0_ADC0_DSUB25_OUT_DQ
PEM (EX)
tatsuki.washimi - 11:54 Wednesday 26 November 2025 (35672) Print this report
Comment to PEM injection 251125 (35662)

I analyzed the PEM injection data:

1. Acoustic injection to PSL room
11/25 12:31:00 - 13:05:00 (JST)
EXC : K1:PEM-EXCITATION_MCF0_RACK_13_EXC
REF : K1:PEM-MIC_PSL_TABLE_PSL1_Z_OUT_DQ
Info : PSL 50 - 900 Hz 171 point 5 Hz resolution 100 cnt excitation

 

The projected noise was dominant around 400Hz widely.

Images attached to this comment
MIF (General)
takaaki.yokozawa - 9:49 Wednesday 26 November 2025 (35671) Print this report
Comment to 'INCREASING LAS POWER' stopped working properly (35668)
1. Check the performance of the HWP
when we rotate the IFO_HWP, the DC PD output of the refl table is decreasing, so IFO HWP worked well.
Also, we performed only the PMC locked and rotate the PSL HWP, the increase of the PSL output increased 0.3 W as expected.
So the HWP would be working well

When I tried to check with IRX again, the behavior was changed, we can increase the laser power without any problem.

When I tried to increase the laser power with PRFPMI RF locked again, the same situation happened again, so we moved to check the UGF of the CARM loop.
Images attached to this comment
MIF (General)
takaaki.yokozawa - 8:45 Wednesday 26 November 2025 (35670) Print this report
Comment to 'INCREASING LAS POWER' stopped working properly (35668)
I also check the state.

The soon after rotate the HWP PSL, locked loss happened as shown in Fig.1. (During the increase las power)
But, in case of the DOWN state (Just locked the IMC), we can increase the laser power as shown in Fig.2.

And in case of the IRX locked state, when we rotate the HWP, the locked loss happened with the similar situation with Fig..1 (Fig.3.)

So maybe some different would be there around the CARM CMS of input 1
Images attached to this comment
MIF (Noise Budget)
kentaro.komori - 2:12 Wednesday 26 November 2025 (35669) Print this report
Measurement of transfer functions and dark noises of IMC CMS

[Fujimoto, Tanaka, Komori]

Abstract:

We measured the transfer functions and dark noises of the IMC common-mode servo (CMS).
The input-referred noise of the IMC CMS seems reasonable, and we found several issues to be resolved in the measurement.

Details:

As a first step toward performing the full CARM noise budget, we measured the transfer functions and dark noises of the IMC CMS using settings as close as possible to those during observation mode.
The power spectra were taken with Mokulab in two frequency ranges: 10 Hz–1 kHz and 1 kHz–100 kHz.
We first measured the dark noise of Mokulab itself (Fig. 1) and compared it with the dark noise of the IMC CMS.

For the measurement of the dark noise of the IMC CMS slow path in the 1 kHz–100 kHz range, we inserted an SR560 because the noise was otherwise hidden beneath the Mokulab noise floor due to the generic filter with a DC gain of 0.1.
The SR560 was configured with a first-order high-pass filter at 300 Hz and a gain of 100 above that frequency.
The input-referred noise of the SR560 is negligible and is also plotted in Fig. 1.
The measured dark noises of the CMS channels are compared as well.

Ideally, the dark noise should be measured with the three boost filters in the common path enabled to fully reproduce the real operating condition.
However, enabling the boost filters caused circuit saturation, so we instead measured the noise without the boost filters and will compensate for their effect later.

The measured transfer functions of the slow and fast paths are shown in Fig. 2.
The corresponding input equivalent noise spectra for both paths are shown in Fig. 3.

Our next steps will be:

  • Resolving the mismatch between the spectra measured in the two different frequency bands.

  • Resolving the discrepancy between the measured and modeled transfer functions.

  • Performing similar measurements for the CARM CMS.

  • Measuring the power dependence of the RFPD noise floor to experimentally determine the shot noise equivalent power.

Images attached to this report
MIF (General)
kentaro.komori - 19:29 Tuesday 25 November 2025 (35668) Print this report
'INCREASING LAS POWER' stopped working properly

After today's works, 'INCREASING LAS POWER' state stopped working properly.
The laser power does not increase in this state now, and we do not understand the reason at this moment.
I requested PRFPMI RF LOCKED state.

Comments to this report:
takaaki.yokozawa - 8:45 Wednesday 26 November 2025 (35670) Print this report
I also check the state.

The soon after rotate the HWP PSL, locked loss happened as shown in Fig.1. (During the increase las power)
But, in case of the DOWN state (Just locked the IMC), we can increase the laser power as shown in Fig.2.

And in case of the IRX locked state, when we rotate the HWP, the locked loss happened with the similar situation with Fig..1 (Fig.3.)

So maybe some different would be there around the CARM CMS of input 1
Images attached to this comment
takaaki.yokozawa - 9:49 Wednesday 26 November 2025 (35671) Print this report
1. Check the performance of the HWP
when we rotate the IFO_HWP, the DC PD output of the refl table is decreasing, so IFO HWP worked well.
Also, we performed only the PMC locked and rotate the PSL HWP, the increase of the PSL output increased 0.3 W as expected.
So the HWP would be working well

When I tried to check with IRX again, the behavior was changed, we can increase the laser power without any problem.

When I tried to increase the laser power with PRFPMI RF locked again, the same situation happened again, so we moved to check the UGF of the CARM loop.
Images attached to this comment
PEM (Tool)
tatsuki.washimi - 17:42 Tuesday 25 November 2025 (35667) Print this report
Rain collector data at Atotsu looks strange

I found the rain collector data at Atotsu looks strange (too small) since June or July this year.
Other channels of the weather station (e.g., temperature) are not strange.

So I suspect some obstacle material are stacked.

Images attached to this report
Comments to this report:
tatsuki.washimi - 16:16 Wednesday 26 November 2025 (35678) Print this report

I checked the rain collector of the Atotsu weather station. It was fullfield by today's rain.

I and Sawada-san cleand it (removing algae).

Images attached to this comment
VAC (Valves & Pumps)
nobuhiro.kimura - 16:38 Tuesday 25 November 2025 (35666) Print this report
Comment to GVbsx, GVbsy were closed suddenly (35640)

[Kimura and M. Takahashi]

 We inspected all vacuum pump units witout the V-chamber areas.
No abnormalities were confirmed in the vacuum pump units of the central mirror room, Y-arm, or Y-end mirror room.
However, the following abnormalities were confirmed:
1. Mortar spray peeling from the ceiling near 530m on the X-arm
2. Looseness in two jacks on the X-10 ion pump support frame
3. Looseness in GVex anchor bolts
Tightening work was performed for items 2 and 3.
Photos are attached for reference.

Images attached to this comment
PEM (Center)
takaaki.yokozawa - 16:06 Tuesday 25 November 2025 (35665) Print this report
Status of the nearby earthquake 251123
Since there are confusion information between earthquake itself and vacuum problem, I set the new port for investigating the earthquake itself.

This earthquake detected at the 18:53:38(JST) 23rd Nov. 2025, that "may" affect to the vacuum level.

Fig.1. showed the
High frequency ground motion (>100 Hz) detected in accelerometer and seismometer

Fig.2. and Fig.3. showed the
Time series for this earthquake.

IXV and EYV arrived at the similar timing, EXV delayed (Expected)
Clear p-wave and s-wave can be seen
There are higher frequency component in X/Y direction compared with Z direction.

Fig.4. and Fig.5. showed the more detail spectrum for this earthquake.
1 Hz resolution, 50 % overlap and 10 times average

Fig.4. IXV/EXV/EYV X/Y/Z axis spectrum with earthquake(red) and without earthquake(blue)
Fig.5. (Left)Comparison of position (IXV(red)/EXV(blue)/EYV(green)) for each axis (X/Y/Z)
Fig.5. (Right) Comparison of axis (X(red)/Y(blue)/Z(green)) for each position (IXV/EXY/EYV)

Now I am trying to make the T-F map by spectgrogram, whitened-spectrogram, Q-transform
Images attached to this report
CAL (General)
hirotaka.yuzurihara - 15:43 Tuesday 25 November 2025 (35661) Print this report
Comment to TCam photo session 20251125 (35660)

The previous photo session was done on 11/14 (klog). So, this shift happened between 11/14 and 11/25.
I heard from Kenta-san, when he performed the initial alighment on 11/21(Fri.), he thought the mirror reference position of ITMY has already shifted. So, I'm not sure this shift is related to the earthquake close to the site (klog).

CAL (Pcal general)
dan.chen - 14:39 Tuesday 25 November 2025 (35664) Print this report
WSK calibration at UToyama

Date: 2025/11/25

Member: Dan Chen, Shingo Hido

We performed our usual WSK calibration at UToyama.

The results look no problem.

Results

Case Alpha (Main Value) Alpha (Uncertainty)
Front WSK, Back GSK -0.911422 0.000068
Front GSK, Back WSK -0.909009 0.000074

Comparison with Previous Results

Comparing with previous results, no significant issues were found.
Attached graph is the result summary including the latest measured data.

Images attached to this report
PEM (EX)
takaaki.yokozawa - 13:56 Tuesday 25 November 2025 (35662) Print this report
PEM injection 251125
I performed the PEM injection tests

1. Acoustic injection to PSL room
11/25 12:31:00 - 13:05:00 (JST)
EXC : K1:PEM-EXCITATION_MCF0_RACK_13_EXC
REF : K1:PEM-MIC_PSL_TABLE_PSL1_Z_OUT_DQ
Info : PSL 50 - 900 Hz 171 point 5 Hz resolution 100 cnt excitation

2. Shake injection to TMSX table EXC1 (Fig.3.)
11/25 13:07:00 - 13:41:00 (JST)
EXC : K1:PEM-EXCITATION_EX0_RACK_1_EXC
REF : K1:PEM-POTABLE_EXC_RACK_EX0_ADC0_DSUB25_OUT_DQ
info : TMSX EXC1 50 - 900 Jz 171 point 5 Hz resolution 100cnt excitation

3. Shaker injection to TMSX table EXC1 sweep
11/25 13:42:30 - 13:47:30 (JST)
EXC : K1:PEM-EXCITATION_EX0_RACK_1_EXC
REF : K1:PEM-POTABLE_EXC_RACK_EX0_ADC0_DSUB25_OUT_DQ
Info : TMSX EXC1 900 -> 50 Hz sweep 100 cnt with 300 sec using awggui

13:47:30 - Silent run
Images attached to this report
Comments to this report:
tatsuki.washimi - 11:54 Wednesday 26 November 2025 (35672) Print this report

I analyzed the PEM injection data:

1. Acoustic injection to PSL room
11/25 12:31:00 - 13:05:00 (JST)
EXC : K1:PEM-EXCITATION_MCF0_RACK_13_EXC
REF : K1:PEM-MIC_PSL_TABLE_PSL1_Z_OUT_DQ
Info : PSL 50 - 900 Hz 171 point 5 Hz resolution 100 cnt excitation

 

The projected noise was dominant around 400Hz widely.

Images attached to this comment
tatsuki.washimi - 12:17 Wednesday 26 November 2025 (35673) Print this report

The channel name in the klog35662 was wrong

wrong K1:PEM-POTABLE_EXC_RACK_EX0_ADC0_DSUB25_OUT_DQ
correct K1:PEM-PORTABLE_EXC_RACK_EX0_ADC0_DSUB25_OUT_DQ
tatsuki.washimi - 12:57 Wednesday 26 November 2025 (35675) Print this report

I analyzed the PEM injection data:

2. Shake injection to TMSX table EXC1 (Fig.3.)
11/25 13:07:00 - 13:41:00 (JST)
EXC : K1:PEM-EXCITATION_EX0_RACK_1_EXC
REF : K1:PEM-POTABLE_EXC_RACK_EX0_ADC0_DSUB25_OUT_DQ
info : TMSX EXC1 50 - 900 Jz 171 point 5 Hz resolution 100cnt excitation

Note that the REF channel name was wrong.

Since a significant excess was not found for almost all frequency injection, the upper limit based on the Coupling Function model (w/o frequency-conversion) is also plotted. 

Images attached to this comment
VAC (Valves & Pumps)
takaaki.yokozawa - 13:55 Tuesday 25 November 2025 (35663) Print this report
Comment to GVbsx, GVbsy were closed suddenly (35640)
Time series for this earthquake.

IXV and EYV arrived at the similar timing, EXV delayed (Expected)
Clear p-wave and s-wave can be seen
There are higher frequency component in X/Y direction compared with Z direction
Images attached to this comment
Search Help
×

Warning

×